OpenStudy (anonymous):

Use an inequality and the five-step process to solve each problem.......Nutrition. Following the guidelines of the U.S. Department of Agriculture, Dale tries to eat at least 5 half-cup servings of vegetables each day. For the first six days of one week, she had 4, 6, 7, 4, 6, and 4 servings. How many servings of vegetables should Dale eat on Saturday, in order to average at least 5 servings per day for the week?

OpenStudy (amistre64):

5step process? that appears to be rather author specific in nature ..

OpenStudy (anonymous):

\[\frac{4+6+7+4+6+4+x}{7}\geq 5\]

OpenStudy (anonymous):

5 step process? like are you making cheese?

OpenStudy (anonymous):

probably learned about 5 step process in math ed course.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

step 1, add the numbers

OpenStudy (amistre64):

these online programs like to make things utterly "specific" in order to determine that you actually read the material they present i believe

OpenStudy (anonymous):

\[\frac{31+x}{7}\geq 5\]

OpenStudy (anonymous):

step 2 multiply by 7 to get \[x+31\geq 35\]

OpenStudy (amistre64):

the downside to that is that they tend to teach a process rather than an understanding .. in my opinion

OpenStudy (anonymous):

step 3 subtract 31 to get \[x\geq 4\]

OpenStudy (anonymous):

step 4 check my arithmetic because the sensei is kicking in

OpenStudy (anonymous):

step 5 ... is that the one where you apologize to everyone, or is that step 8?

OpenStudy (amistre64):

lol .... the 12 step program for mathaholics ;)

OpenStudy (anonymous):

i like it

OpenStudy (anonymous):

no joking please if i dont get this problem im going to fail.

OpenStudy (amistre64):

it would help if we knew what your 5 steps HAVE to be ...

OpenStudy (amistre64):

the math here doesnt take any preset 5 steps that are universally accepted.

OpenStudy (amistre64):

the concept tho is to understand what it means to get an average; and an average is simply adding up all the values and dividing that amount by how many values there are

OpenStudy (anonymous):

forget amistre64., i give up//

OpenStudy (amistre64):

hmm, there is no reason to give up. Its just that in order to give an adequate answer to the question there is information that we simply do not have. And that would be the requirement of "5 steps". But in the end, good luck with it :)

OpenStudy (anonymous):

what information do you need

OpenStudy (amistre64):

the "5 steps" information which your online material covers and asks you to use. That information is content specific, and not a universal mathematical concept.